### v2

# Campaign-spending driven polarization transition in a double-random field model of elections

Abstract
We model bipartisan elections where voters are exposed to two forces: local homophilic interactions and external influence from two political campaigns. The model is mathematically equivalent to the random field Ising model with a bimodal field. When both parties exceed a critical campaign spending, the system undergoes a phase transition to a highly polarized state where homophilic influence becomes negligible, and election outcomes mirror the proportion of voters aligned with each campaign, independent of total spending. The model predicts a hysteresis region, where the election results are not determined by campaign spending but by incumbency. Calibrating the model with historical data from US House elections between 1980 and 2020, we find the critical campaign spending to be $\sim 1.8$ million USD. Campaigns exceeding critical expenditures increased in 2018 and 2020, suggesting a boost in political polarization.
Classical statistical physics provides a framework for understanding collective phenomena. It typically assumes weakly interacting constituents, governed by a single, time-invariant force, which allows computation of phase diagrams and identification of critical points where system-wide change occurs. This framework has long been applied to social phenomena, including financial markets [1, 2, 3], pedestrian and crowd dynamics [4, 5], anthropology [6, 7], judicial systems [8], and opinion formation [9, 10, 11, 12]. In particular, it underpins opinion dynamics models such as bounded confidence [13, 14], social validation [15], and cultural dissemination [16].
In voting behavior, one of the most influential physics-inspired models is the voter model [17], where agents adopt neighbors’ opinions on lattices or heterogeneous networks [18]. Extensions include stochastic noise [19], committed agents or “zealots” [20], and the interplay of micro- and macrodynamics [21]. It has been shown that the voter model reproduces statistical features of US presidential elections [22]; see also [23] for a review.
A more recent class of opinion dynamics models builds on two sociological principles: homophily—the tendency to associate with similar others [24]—and social balance—the reduction of cognitive dissonance in triads [25, 26]. Extending the voter model, these approaches capture more realistic interactions. Their combined effects have been studied jointly [27, 28] and integrated into a unified framework [29], applied to group formation [30] and recently validated experimentally [31]. Many such models, including voter- and homophily-based ones, are inspired by the Ising model, long a versatile tool in interdisciplinary contexts [32]. In elections, it explained the universal scaling of vote distributions in proportional systems [33], predicted margins of victory from turnout [34], and modeled temporal effects through an external field [35].
A central challenge in complex adaptive systems such as societies is the variety of time-dependent interactions, co-evolution, herding, and anticipation of others’ actions [36]. While some aspects are easy to model, calibration and validation remain difficult, requiring great care to ensure testable models. Here, we address opinion formation with multiple interaction types, as in political elections. Voters exchange views within social networks of family, friends, and colleagues, while also following political campaigns—typically only those of their preferred party. We model these two processes through Ising interactions for homophily and a bimodal random field for campaign influence. Together they form a Random Field Ising Model (RFIM), extending spin-spin interactions with site-dependent random external fields.
Despite its simplicity, the RFIM captures rich behavior such as quenched disorder and complex phase diagrams [37, 38, 39]. Variants with bimodal random fields, where the field takes two values [40, 41, 42], exhibit tricritical points marking the transition from second- to first-order phase changes [43, 44]. In our context, the RFIM represents a bipartisan electorate where each voter holds a binary preference. Voters occupy a social (friendship) network and are randomly assigned one of two field values: the sign encodes campaign affiliation, while the magnitude reflects campaign strength, with spending serving as a proxy. A schematic illustration is shown in End Matter, Fig. 3.
The aim of this paper is to study the interplay between homophily and campaign-following the emergence of campaign polarization, defined as the normalized difference between average opinions of voters exposed to each campaign. Low polarization indicates that both groups vote similarly, with decisions mainly shaped by homophily, whereas high polarization means groups align with campaigns and are less influenced by neighbors. We compute the phase diagram as a function of the “temperature,” representing susceptibility to opinion change, and the campaign spending of the two parties.
We focus on the effects of increasing campaign intensity. At low spending, opinions are shaped mainly by homophily, but as campaign influence grows, voter preferences are increasingly influenced by campaign messaging. Key questions are: when does campaign alignment outweigh homophilic similarity, how does polarization evolve at this point, and how does this transition affect social tension and election outcomes? A strength of our model is that it can be calibrated and tested on empirical data; we use particularly the US House elections between 1980 and 2020. Calibration allows us to infer the “temperature” and the critical spending threshold above which campaign polarization rises sharply, and to track how many races exceeded this threshold over four decades.
Election model as an RFIM with a bimodal field. — We consider $N$ voters with binary opinions $s_{i}\in{\pm 1}$, representing preference for one of two parties in a bipartisan election. Voters form a social network encoded in the adjacency matrix $A_{ij}$ and interact through homophily, tending to align with neighbors. Each voter also follows one of the two campaigns, modeled by an external field $h_{i}$ drawn from a bimodal distribution,
$p(h_{i})=p\,\delta(h_{i}-h^{+})+(1-p)\,\delta(h_{i}+h^{-}),$ (1)
where $\delta(x)$ is the Dirac delta function and $p\equiv p(h^{+})$ is the probability of following the first party’s campaign. With $h^{+},h^{-}\geq 0$, the field takes values $h^{+}$ or $h^{-}$. The system Hamiltonian is
$H(s_{1},\dots,s_{N})=-J\sum_{i<j}A_{ij}s_{i}s_{j}-\sum_{i}h_{i}s_{i}.$ (2)
To solve the model, we apply two approximations: the configuration model and a mean-field approximation (see Supplemental Material). Denoting the average magnetization by $m=\langle s_{i}\rangle$, we arrive at the mean-field Hamiltonian $H^{MF}(s_{1},\dots,s_{N})=-\sum_{i}(\tilde{J}m+h_{i})s_{i}$, where $\tilde{J}=J\langle k\rangle$. The equilibrium distribution is therefore $p(s|h^{\pm})=\exp\left[-\beta(\tilde{J}m\pm h^{\pm})s\right]/Z^{\pm}$, where $\beta=(kT)^{-1}$ is the inverse temperature (for the rest of the paper, we set $k=1$), and $Z^{\pm}=2\cosh\left[\beta(\tilde{J}m\pm h^{\pm})\right]$ is the partition function. Here, the temperature $T$ represents social volatility—the willingness of individuals to adopt new opinions, even if this increases social stress.
The average magnetization under field $\pm h^{\pm}$ is $m^{\pm}\equiv\langle s\rangle^{\pm}=\tanh[\beta(\tilde{J}m\pm h^{\pm})]$. The population magnetization is $m=pm^{+}+(1-p)m^{-}$, representing the election outcome ($m=\pm 1$ corresponds to an unanimous result; $m=0$ represents 50:50 split). It satisfies the self-consistency equation,
$m=p\tanh[\beta(\tilde{J}m+h^{+})]+(1-p)\tanh[\beta(\tilde{J}m-h^{-})]\,.$ (3)
Setting $m=0$ yields
$p\tanh(\beta h^{+})=(1-p)\tanh(\beta h^{-})\,$, (4)
which reduces to $h^{+}=h^{-}$ for $p=\tfrac{1}{2}$. We define campaign polarization as $\pi=\tfrac{1}{2}(m^{+}-m^{-})$, the difference between average opinions of voters following opposite campaigns. If both groups share the same opinion, $\pi=0$; if they hold opposite views ($m^{+}=1$, $m^{-}=-1$), then $\pi=1$.
Critical parameters. — We first summarize the known results for the symmetric case $p=\tfrac{1}{2}$ and $h^{+}=h^{-}\equiv h$. We set $\tilde{J}=1$. As shown in [44] and in the Supplemental Material, the model exhibits a continuous crossover for $T>1$. For $T<1$, it undergoes a second-order transition at $h_{c}=T\,\textrm{arctanh}\left(\sqrt{1-T}\right)$. A first-order transition occurs for lower temperatures, with tricritical point $T^{*}=\tfrac{2}{3}$ and $h^{*}=\tfrac{2}{3}\,\textrm{arctanh}\left(\tfrac{1}{\sqrt{3}}\right)\approx 0.439$.
We extend this result to the non-symmetric case using Eq. (4), which links $h^{+}$ and $h^{-}$, and by expanding the self-consistency equation (3) around $m=0$. Unlike the symmetric case, the quadratic term does not vanish, yielding the critical curves
$h_{c}^{+}=T\,\textrm{arctanh}\left(\sqrt{(1-T)\frac{1-p}{p}}\right)\,$, (5)
$h_{c}^{-}=T\,\textrm{arctanh}\left(\sqrt{(1-T)\frac{p}{1-p}}\right)\,.$ (6)
The full derivation is given in the Supplemental Material.
Figure 1: Phase diagrams of the election model. Phase diagram for magnetization, $m$ (a-d), and polarization, $\pi$, (e-h) in the $(h^{+},h^{-})$ plane for temperature $T=1$ (a,c,e,g), and for $T=0.75$ (b,d,f,h), and a prior probability, $p=0.5$ (a,b,e,f) and $p=0.6$ (c,d,g,h). The black dashed line shows $m=0$. The purple point marks the maximal point of the hysteresis, as derived in the main text. For $T=1$, the expected behavior is that the magnetization is directly affected by the relative strength of the two fields, affected by pp. For $T=0.75$, we observe a more interesting behavior of the phase diagram. For the case of low field strength, we observe hysteresis (striped region). In both cases, the campaign polarization, $\pi$, starts to increase rapidly when both field strengths are above a critical value, $h_{c}$ (red).
Phase diagram. — In the $(h^{+},h^{-})$ plane, we solve Eq. (3) numerically to obtain the phase diagrams shown in Fig. 1. For $T\geq 1$, the system has a single stable solution: the candidate with higher campaign spending wins, with the boundary given by Eq. (4). For $T<1$ and low fields, a hysteresis region appears around the curve given by Eq. (4) until reaching the critical values of Eqs. (5) and (6). This implies voter behavior depends on prior states, which we interpret as an incumbency effect where officeholders retain an advantage even with lower spending than predicted by Eq. (4). The incumbency effect is well documented [45, 46] and considered central to campaign strategy [47].
Campaign polarization $\pi$ remains near zero when at least one external field is weak and increases only when both exceed their critical values, $h^{+}\gtrsim h^{+}_{c}$ and $h^{-}\gtrsim h^{-}_{c}$. Like magnetization, polarization undergoes a phase transition for $T<1$. In the high-polarization regime, the overall magnetization is nearly constant, $m\approx 2p-1$, as predicted by Eq. (4) for $h^{+},h^{-}\gg 1$, an effect most pronounced at low $T$. Thus, campaign influence dominates homophily: voters aligned with a campaign tend to vote uniformly for that party, regardless of neighbors. This matches recent results [48] linking strong campaign polarization to voter extremism, where voters adopt increasingly extreme positions under intense pressure. Finally, polarization here corresponds to affective polarization [49, 50, 51], where individuals bond more with their political group than with ideology. Our model captures this as the interplay of ideological and partisan homophily, amplified by campaign intensity.
Refer to captionRefer to caption
Figure 2: Estimation of model parameters for US House of Representatives. We compare campaign spending and election results for 6357 of 9135 races between 1980–2020, focusing on close races ($p=0.5\pm 0.05$). (a) For $T\geq 1$, the classification model (see Supplemental Material) predicts the higher-spending candidate wins. (b) For $T<1$, it predicts an incumbency region (yellow) where incumbents win despite lower spending. (c) Optimal parameters $T$ and $h_{c}$ are estimated by maximizing classification accuracy across all 6357 races, yielding $T^{=}0.922$ and $h_{c}=\mathdollar 1.83$M. Cases where incumbents win with lower spending are highlighted, with additional black borders for points in the hysteresis region. The spending diagram is truncated to show the incumbency region (yellow). (d) Accuracy across $T$ is shown, with the maximum marked by a red star. The inset shows a McNemar contingency table comparing the optimal model ($T=T^{*}$) to the null model ($T=1$). The McNemar test gives $p<0.0001$, indicating significantly better performance of the optimal model.
Calibration to US House election data. — To test the model against real data, we analyze all US House campaigns in 435 districts across 21 elections (presidential and midterms) from 1980–2020. Campaign spending and results are publicly available via the Federal Election Commission (FEC) [52] and in machine-readable form at [53]. We focus on House races because congressional districts have relatively uniform populations, enabling meaningful spending comparisons. To ensure bipartisan competition, we restrict to races contested solely between Democratic and Republican candidates, excluding those with significant third-party or independent contenders. This yields 6357 races from 9135 in the period. All campaign expenditures are inflation-adjusted to 2020 USD using the Consumer Price Index (CPI). For each race, the campaign share parameter $p$ is set by the previous election result in the same district. For the first election in our dataset (1980), $p$ is taken from the 1978 publicly available results.
To estimate model parameters, we build a classification framework based on the proposed dynamics to predict race winners. Outside the hysteresis region, the winner is determined by the sign of total magnetization; within the hysteresis region, outcomes depend on incumbency when an incumbent is present. This scheme is illustrated in Fig. 2(a,b) and detailed in the Supplemental Material.
We first consider the symmetric case $p\approx 0.5$. For $p=0.5$, the classification model predicts the same outcome for all $T\geq 1$: the higher-spending candidate wins (Fig. 2(a)). For $T<1$, a hysteresis region emerges (yellow) where incumbency dominates. This region centers near $h^{DEM}=h^{REP}$, with its shape depending on $T$ and $h_{c}$ (Fig. 2(b)). By fitting the empirical boundaries of this region, we infer $T$ and $h_{c}$. Optimal parameters are estimated by maximizing classification accuracy over $T$ and $h_{c}$, yielding $T^{*}=0.922$ and $h^{*}=\mathdollar 1.83M$ (Fig. 2(c)). The figure highlights all closely contested races ($p=0.5\pm 0.05$), marking cases where incumbents win despite lower spending (black-bordered dots).
In the End Matter, we present a similar analysis for races with $p\approx 0.6$, i.e., Republican-leaning districts ($p=0.6\pm 0.05$). Applied to this subset, the parameters shift slightly ($T=0.845$, $h^{*}\sim\mathdollar 2M$), but the qualitative behavior — hysteresis and emergence of polarization — remains unchanged.
Figure 2(d) shows classification accuracy as a function of temperature, with $h_{c}$ chosen at each $T$ to maximize accuracy. To test statistical significance against a baseline, we compare the optimal classifier to a null model without hysteresis ($T\geq 1$) using the McNemar test (see Supplemental Material). The inset shows the contingency table of correctly and incorrectly classified results for both models.
We find a statistically significant improvement in accuracy ($\chi^{2}=24.69$, p-value $p=6.76\cdot 10^{-7}$), though the absolute gain is small. This reflects that in most races incumbents outspend challengers, so both the null model ($T\geq 1$) and optimal model ($T=T^{\star}$) yield similar predictions. This aligns with prior findings that incumbency strengthens fundraising, with incumbents typically attracting more support [54].
In the End Matter, we further extend our analysis to the emergence of polarization. By estimating the critical spending threshold $h_{c}$, we test how many races fall in the polarization region, i.e., when both parties exceed $h_{c}$ (Fig. 4). Panel (a) shows the full spending region including the polarized area, while panel (b) presents results for $p\approx 0.5$. At low spending, outcomes are decisive; above $h_{c}$, results cluster near 50:50, consistent with the prediction that for $p\approx 0.5$ most outcomes in the polarized region satisfy $m\approx 0$. Panel (c) compares the number of races where both campaigns exceeded $h_{c}$ across election cycles, revealing a sharp rise in 2018 and 2020. While these results rest on model predictions and need further empirical validation, the trend matches recent observations of rising campaign polarization [55].
Discussion. — We introduced a simple election model combining two mechanisms influencing voter decisions: homophily, i.e., interactions with family, friends, and close contacts, and political campaign influence. Despite its simplicity, the model can be calibrated with real-world data (US House elections) and reproduces rich behavior absent in earlier work. The fact that voters typically follow only one campaign leads to a phase transition: campaign polarization rises sharply once both parties exceed a critical spending threshold. In this regime, most outcomes are 50:50 when $p\approx 0.5$ (swing states), regardless of the detailed spending. For biased states ($p\neq 0.5$), the model allows us to estimate the challenger’s minimum spending above which 50:50 outcomes become possible.
We identified a hysteresis region in the phase diagram that leads to an incumbency effect, enabling quantitative assessment of incumbents’ advantage. The model shows that challengers must overcome an initial threshold of about $\mathdollar 140,000$ even if incumbents spend nothing. This barrier decreases with higher overall spending but remains significant: for instance, when the incumbent spends $0.5h_{c}$ ($\approx\mathdollar 900,000$), the challenger still faces a disadvantage of about 20% of total campaign cost. This quantifies a structural incumbency advantage beyond candidate-specific factors. By quantifying incumbency strength, we estimate the effective temperature of the social system and identify a polarization threshold. Notably, even without a universal definition of “temperature” in social contexts, it can be inferred indirectly from observable critical phenomena such as hysteresis and field-driven phase transitions.
Several studies reported a decline in incumbency strength in recent decades [56], raising the question of whether this relates to shifts in effective “temperature”. It remains open whether this temperature is universal or varies across elections, contexts, or regions. Since incumbency is central to electoral strategy [57], further work should examine how decisions such as planned retirements shape challengers’ prospects by mitigating incumbents’ inherent advantage.
The model can be naturally extended in several ways. Multipartisan systems can be described by an extension to a double-random field Potts model [58] with more than two opinion states. More realistic scenarios could be added through extensions to heterogeneous friendship networks, explicit party membership, or mechanisms such as primaries. Homophily may also vary across ties—for instance, individuals might ignore co-workers’ views but adopt those of parents or close friends. Such modifications would better capture electoral complexity. Moreover, campaign intensity may not scale linearly with spending, as noted in earlier work [59, 60], suggesting future models should consider nonlinear or context-dependent effects.
Despite its simplicity, the model predicts complex game-theoretic behavior. Strategic aspects of campaigns have been studied from this perspective [61, 62], and our results deepen this understanding. The model suggests that candidates may rationally raise spending to induce polarization; once reached, it is difficult for opponents to reverse, locking in an advantage. This is especially effective in partisan districts, where dominant candidates benefit from driving the electorate into a polarized phase. Yet this strategy carries social costs: polarization erodes ties across divides and exacerbates fragmentation. A key implication of the model is that regulatory interventions, such as caps on campaign spending, may be needed to prevent undesirable outcomes. While spending limits are under renewed discussion [63], our results add a novel perspective by emphasizing the broader societal costs of unbounded campaign intensity.
We shall also note that similar results were found in the context of polarization in the US Congress itself [64]. The study shows that the polarization in the US Congress increased after the 2010 Supreme Court approval of Super PACs, which enabled an increase in donor influence.
Elections are not the only context where homophily competes with external influence. Similar dynamics appear in marketing, where peer effects and advertising shape consumer behavior. A classic case is the Coca-Cola vs. Pepsi rivalry [65], showing how competing campaigns polarize preferences. Our framework illustrates how polarization generally arises from the tension between social ties and persuasive efforts.
Acknowledgments. — We thank Mirta Galesic and Henrik Olsson for helpful discussions. We acknowledge support from the Austrian Science Fund (FWF) project No. P 34994 (grant-DOI: 10.55776/P34994), the Austrian Research Promotion Agency (FFG) as part of the ESSENCSE project (873927), and the Austrian Federal Ministry for Climate Action, Environment, Mobility, Innovation & Technology (BMK) as part of the CSH Postdoc Programme (GZ 2023-0.841.266).
Supplemental Material
Detailed derivation of the self-consistency equation. We start with the Hamiltonian
$H(s_{1},\dots,s_{N})=-J\sum_{i<j}A_{ij}s_{i}s_{j}-\sum_{i}h_{i}s_{i}\,$,
In order to decouple the Hamiltonian into the Hamiltonian for a single individual, we use the two approximations. The first one is the configuration model approximation, where we assume that the adjacency matrix of a random network can be approximated $A_{ij}\approx\frac{k_{i}k_{j}}{Nk}$ where $k_{i}$ is the degree (i.e., connectivity) of the node $i$ and $\langle k\rangle$ is the average degree. The second approximation we use is the mean-field approximation. Expressing the spin in terms of its average value as $s_{i}=\langle s_{i}\rangle+\delta s_{i}$ enables us to omit the term quadratic in fluctuations. By denoting the average magnetization as $m=\langle s_{i}\rangle$, we can rewrite the Hamiltonian as
$H(s_{1},\dots,s_{N})$ $\approx$ $\frac{Jm^{2}N\langle k\rangle}{2}-J\sum_{i}\langle k\rangle ms_{i}-\sum_{i}h_{i}s_{i}$
The first term can be omitted from the Hamiltonian since it is just an additive constant to the energy, which will be canceled out once the equilibrium distribution is calculated. Thus, the mean-field Hamiltonian can be expressed as
$H^{MF}(s_{1},\dots,s_{N})=-\sum_{i}(\tilde{J}m+h_{i})s_{i}$
where $\tilde{J}=J\langle k\rangle$. Note that $h_{i}$ is a random variable with the distribution
$p(h_{i})=p\delta(h_{i}-h^{+})+(1-p)\delta(h_{i}+h^{-})\,$,
The equilibrium distribution is therefore
$p(s|h^{\pm})=\exp\left(-\beta(\tilde{J}m\pm h^{\pm})s\right)/Z^{\pm}$
where $\beta=\frac{1}{kT}$ is the inverse temperature, and
$Z^{\pm}=2\cosh\left(\beta(\tilde{J}m\pm h^{\pm})\right)$
is the partition function. The average magnetization of spins coupled to external field $\pm h^{\pm}$ is
$m^{\pm}=\langle s\rangle^{\pm}=\sum_{s=\pm 1}sp(s|h^{\pm})=\tanh\left(\beta(\tilde{J}m\pm h^{\pm})\right)$
. The total magnetization can be expressed as
$m$ $=$ $\langle m\rangle_{h}=pm^{+}+(1-p)m^{-}$
$=$ $p\tanh\left(\beta(\tilde{J}m+h^{+})\right)+(1-p)\tanh\left(\beta(\tilde{J}m-h^{-})\right)\,.$
Derivation of the critical curve and tricritical point for symmetric case. — Let us now focus on the symmetric case, i.e., when $p=\frac{1}{2}$ and $h^{+}=h^{-}\equiv h$. We investigate how the phase diagram depends on the strength of the field and the temperature. Without loss of generality, we consider that $\tilde{J}=1$. To this end, we expand the right-hand side of the self-consistency equation around $m=0$ and get
$m=\frac{1-\xi^{2}}{T}m+\frac{4\xi^{2}-3\xi^{4}-1}{3T^{3}}m^{3}+\mathcal{O}(m^{5})$
where $\xi\equiv\tanh(h/T)$. This cubic equation has three solutions: one trivial $m^{0}=0$ and two non-trivial solutions
$m^{\pm}=\pm\frac{\sqrt{3(1-T-\xi^{2}})}{\sqrt{\frac{T^{2}+3\xi^{4}-4\xi^{2}}{T^{2}}}}$
By comparing when the non-trivial solutions become trivial, i.e., $m^{0}=m^{\pm}$, we obtain the critical curve
$h_{c}=T\textrm{arctanh}(\sqrt{1-T})$
Finally, by plugging the critical curve into the self-consistency equation, we can determine the order of the phase transition from the sign of the third-order coefficient. The coefficient is along the critical curve equal to $\frac{2-3T}{3T^{2}}$ so the critical point where the phase transition changes its order is
$T^{*}=\frac{2}{3}\quad h^{*}=\frac{2}{3}\textrm{arctanh}\left(\frac{1}{\sqrt{3}}\right)\approx 0.439$
Derivation of the critical curve and tricritical point for the asymmetric case. Let us now focus on the general case when $p\neq\frac{1}{2}$. We take the condition for $m=0$, which is
$p\tanh(\beta h^{+})=(1-p)\tanh(\beta h^{-})$
and expand the self-consistency equation
$m=p\tanh\left(\beta(\tilde{J}m+h^{+})\right)+(1-p)\tanh\left(\beta(\tilde{J}m-h^{-})\right)\,.$
around $m=0$ while keeping the dependence between $h^{+}$ and $h^{-}$ determined from the condition on $m=0$ above.
Let us use the Taylor expansion of $\tanh(\beta(m+h))$ which is
$\tanh(\beta(m\pm h^{\pm}))=\pm\tanh(\beta h^{\pm})$
$+\beta m\left(1-\tanh^{2}(\beta h^{\pm})\right)$
$\mp\beta^{2}m^{2}\tanh(\beta h^{\pm})(1-\tanh^{2}(\beta h^{\pm}))+\mathcal{O}(m^{3})$
By denoting $\xi^{\pm}=\tanh(h^{\pm}/T)$, one can rewrite the self-consistency equation as
$m=p\left[\xi^{+}+m\frac{1-(\xi^{+})^{2}}{T}-m^{2}\frac{\xi^{+}(1-(\xi^{+})^{2})}{T^{2}}\right]$
$+(1-p)\left[-\xi^{-}+m\frac{1-(\xi^{-})^{2}}{T}+m^{2}\frac{\xi^{-}(1-(\xi^{-})^{2})}{T^{2}}\right]+\mathcal{O}(m^{3})$
Since $p\xi^{+}-(1-p)\xi^{-}=0$, the constant term is zero; therefore, we can rearrange the terms as
$\left[\frac{T-p(1-(\xi^{+})^{2})-(1-p)(1-(\xi^{-})^{2})}{T}\right]m$
$+\left[\frac{p\xi^{+}(1-(\xi^{+})^{2})-(1-p)\xi^{-}(1-(\xi^{-})^{2})}{T^{2}}\right]m^{2}=0$
The solution is therefore either $m_{0}=0$ or
$m_{1}=-T\frac{T-p(1-(\xi^{+})^{2})-(1-p)(1-(\xi^{-})^{2})}{p\xi^{+}(1-(\xi^{+})^{2})-(1-p)\xi^{-}(1-(\xi^{-})^{2})}\,.$
Now, the critical point is given by the condition $m_{1}=m_{0}\equiv 0$, which is equivalent to
$T=p(1-(\xi^{+}_{c})^{2})+(1-p)(1-(\xi^{-}_{c})^{2})\,.$
By plugging in from the condition $\xi^{-}=\frac{p}{1-p}\xi^{+}$, we obtain
$T=p(1-(\xi^{+}_{c})^{2})+(1-p)(1-\left(\frac{p}{1-p}\right)^{2}(\xi^{+}_{c})^{2})$
$=1-\frac{p}{1-p}(\xi^{+}_{c})^{2}\,.$
By plugging back for the $\xi_{c}^{+}=\tanh(h^{+}_{c}/T)$, we can express $h_{c}^{+}$ on $T$ as
$h_{c}^{+}=T\textrm{arctanh}\sqrt{(1-T)\frac{1-p}{p}}\,.$
Similarly, by expressing $h_{c}^{-}$ from the condition, we get that
$h_{c}^{-}=T\textrm{arctanh}\sqrt{(1-T)\frac{p}{1-p}}\,.$
Interestingly, for $p=\frac{1}{2}$, both equations boild down to $h_{c}^{+}=h_{c}^{-}\equiv h_{c}$.
Classification model. Here, we describe the classification models used in the main text. The classification model is directly based on the results of the Random Field Ising model. Without loss of generality, we assign $h^{+}=h^{REP}$ as the spending of the Republican party candidate, $h^{-}=h^{DEM}$ as the spending of the Democratic party candidate. The prediction of the classification model, based on the magnetization $m$ goes as follows:
  • If $(h^{DEM},h^{REP})$ lie in the hysteresis region, then the model predicts the incumbent as the winner.
  • If $(h^{DEM},h^{REP})$ lie outside of the hysteresis region, or if there is an open seat (i.e., the incumbent does not candidate), a Republican wins if $m>0$, and a Democrat wins if $m<0$.
Since, for the temperature $T\geq 1$, we observe no hysteresis in the region, only the second condition applies. Specifically, when $p=0.5$, the condition on the sign of magnetization $m$ boils down to the condition whether $h^{REP}>h^{DEM}$ (corresponding to $m>0$) or the other way around. In this case, the classification model does not depend on temperature $T$ (when $T\geq 1)$. We call this model the null model. This model catches the intuitive idea that in the case of equal campaign coverage, the candidate who spends more money on the campaign wins the election.
Model accuracy. In order to measure the performance of the classification model, we use the model accuracy. The confusion table between the predicted classification and the actual classification (here, the prediction is that a Republican candidate wins an election) is then defined as
predicted/actual
classification positive negative
positive True positive (TP) False negative (FN)
negative False positive (FP) True negative (TN)
The accuracy is defined as
$ACC=\frac{n_{TP}+n_{TN}}{n_{TP}+n_{TN}+n_{FP}+n_{FN}}$
McNemar test. The McNemar test is used to demonstrate whether one of the two classification models used on a given data set is better than the other. For each observation, a classification model gives a predicted classification, which is compared with the actual classification. For example, in the election races, the classification model predicts the winner of the election based on the campaign spending and incumbency (see the second above), which is then compared with the actual election result.
For two classification models $M_{1}$$M_{2}$, the contingency table between correctly and incorrectly classified observations can be written as follows:
$M_{1}$$M_{2}$ $M_{2}$ correct $M_{2}$ wrong
$M_{1}$ correct $n_{11}$ $n_{12}$
$M_{1}$ wrong $n_{21}$ $n_{22}$
The null hypothesis is that both marginals are the same, and therefore the probability that the first model is correct and the second model is wrong is the same as that the first model is wrong and the second model is correct. The McNemar test statistic is
$\chi^{2}=\frac{(n_{12}-n_{21})^{2}}{n_{12}+n_{21}}\,.$
Under the assumption of the null hypothesis, and for large enough $n_{12}$ and $n_{21}$, the statistic follows a $\chi^{2}$ distribution with one degree of freedom. We can therefore reject the null hypothesis if the observed statistic is significant (i.e., the $p$-value is smaller than the desired statistical level).
Calibration to US House election data for Republican-leaning races ($p=0.6\pm 0.05)$. To illustrate the effectiveness of the classification model also on the case of a subset of races, we choose the races whose previous results were in the range corresponding to $p=0.6\pm 0.05$. The subset consists of 1145 election races. We find the optimal classification model by finding $T$ and $h_{c}$ that maximize accuracy. We find out that while the optimal parameters slightly change ($T^{\star}=0.845$, $h_{c}=2$ million USD), the overall behavior does not change. We should also stress that since the smaller size of the subset, the statistical tests like the McNemar test exhibit a bit weaker (but still significant) value, which is caused by the fact that the number of points that are classified differently by the optimal model ($T=T^{\star}$) and the null model ($T=1$) is relatively low.